OpenStudy (anonymous):

Write the equation of the parabola that has a vertex (1, -1) and passes through the point (2,5

OpenStudy (jlastino):

the equation of a parabola is: y = a(x-h)² + k where (h,k) is the vertex. since you are given (h,k) and (x,y), you can find "a" by plugging in the know values and solving: 5 = a(2-1)² - 1 5 = a -1 a = 6 a = -5/4 quadratic function: y = f(x) = (6)(x - 1)² -1 6(x^2-2x+1)-1 y=6x^2-12x+5